OpenStudy (anonymous):

A sound wave is modeled with the equation y =1/4 cos 2pi/3 theta a. Find the period. Explain your method. b. Find the amplitude. Explain your method

OpenStudy (phebe):

B XD

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 4 } \cos 2\pi/3\]

OpenStudy (anonymous):

there's supposed to be a theta at the end don't know why it didn't show

OpenStudy (anonymous):

can you help me answer this?

OpenStudy (phebe):

idk im not sure hold on

OpenStudy (phebe):

nm sorry @mathmale help him or her

OpenStudy (anonymous):

ok, I looked at some different answers to the same question and didn't really get how they got the answer. http://openstudy.com/updates/519e5489e4b04449b22151a5

OpenStudy (phebe):

opk hold on

OpenStudy (anonymous):

you know what I actually think I figured it out myself, thank you anyways!